OpenStudy (danielannx3):

Ellen wants to purchase a book that is regularly priced at $18. The book is discounted 15%. She also needs to pay a 6% sales tax on the discounted price. What is the total amount Ellen will pay for the book?

OpenStudy (ilovepuppies):

$18 - 15 + 6 = ?

OpenStudy (danielannx3):

i think so let me try it thank you tho @ILovePuppies

OpenStudy (danielannx3):

can someone help i multiply 0.15*18=2.70 so that give me 18-2.70=15.30 what do i do next

OpenStudy (danielannx3):

@agent0smith

OpenStudy (danielannx3):

@mathmate

OpenStudy (danielannx3):

someone please help

OpenStudy (agent0smith):

If it's discounted 15%, that means it's 85% of it's original price. So find 85% of the original price. Then find 6% of that new discounted price; that's the sales tax. Add it to the discounted price.

OpenStudy (danielannx3):

so she will pay $4.32? @agent0smith

OpenStudy (agent0smith):

Almost definitely not... show your work. And do it in the order I suggested in the last post.

OpenStudy (danielannx3):

0.85*18=15.30 18-15.30=2.70 0.6*2.70=1.62 2.70+1.62=4.32 Thats what i got @agent0smith

OpenStudy (agent0smith):

You're doing some of the right working, but you didn't follow what I said. You did other steps that aren't correct. And 6% is not 0.6.

OpenStudy (agent0smith):

"find 85% of the original price. Then find 6% of that new discounted price; that's the sales tax. Add it to the discounted price."

OpenStudy (queenie):

thats because there is 2 ways to do this....agent0smith is doing it the easiest way... since it is 15% off, she is paying 85% for it. so take 85% of 18...and then add 6% of 18 to that. OR... take 15% of 18...and subtract that from 18....then take 6% of 18..and add that to ur previous amount.